Derive an expression for the hydrolysis constant of sodium acetate.
1
Expert's answer
2019-10-28T07:50:29-0400
"Hydrolysis\ of\ Sodium\ acetate\ -\\CH_3COO^-+ H_2O \iff CH_3COOH+OH^-"

equilibrium constant equation -


"K_e=\frac{[CH_3COOH][OH^-]}{[CH_3COO^-][H_2O]}"

Concentration of water is very large and is regarded as practically constant so it will assumes the

form -

"K_h=\frac{[CH_3COOH][OH^-]}{[CH_3COO^-]}\ \ \ \ .......(1)"

but,


"K_w=[H^+][OH^-]\ \ \ \ ........(2)"

For the dissociation of weak acid acetic acid,


"CH_3COOH\iff H^++CH_3COO^-"

acid dissociation constant, "K_a" is expressed as


"K_a=\frac{[H^+][CH_3COO^-]}{[CH_3COOH]}\ \ \ \ \ .......(3)"

Dividing (2) by (3)


"\frac{K_w}{K_a}=\frac{[OH^-][CH_3COOH]}{[CH_3COO^-]}=K_h"

or,


"K_h=\frac{K_w}{K_a}"